OpenStudy (pokiedokie):

simplify the expression using the imaginary unit i: 2√-24 + 6 simplify the expression: (2-3i)/(1+5i) please help and walk me through the steps!

OpenStudy (lochana):

hint:\[i^2 = -1\]

OpenStudy (lochana):

\[2\sqrt{-1.24} + 6\]

OpenStudy (lochana):

still confused?

OpenStudy (pokiedokie):

do i distribute the 2 and then add 6 or is that not important?

OpenStudy (lochana):

no. you can put i^2 instead of -1\[2\sqrt{i^224} + 6\]

OpenStudy (lochana):

get it?

OpenStudy (lochana):

i is the imaginary unit here.okay?

OpenStudy (pokiedokie):

yes

OpenStudy (lochana):

so we can say\[2\sqrt{i^2}\sqrt{24}+ 6\]

OpenStudy (lochana):

look at my previous one if you don't understand last one

OpenStudy (lochana):

I think you still have problems.. right?

OpenStudy (pokiedokie):

im just taking a while to respond because i'm trying to understand it better

OpenStudy (lochana):

okay. that's fine. do you know why we use this i^2 thing here?

OpenStudy (pokiedokie):

no not really, sorry

OpenStudy (lochana):

because you can't solve \[\sqrt{-24}\]

OpenStudy (lochana):

you can't square root negative numbers.

OpenStudy (pokiedokie):

ohhh, that makes sense, so what it's trying to do is make it positive then

OpenStudy (lochana):

yes. so we get rid of negative part using i^2. we say \[i^2 = -1\]

OpenStudy (pokiedokie):

okay i see so it turns into positive 24 and then we square root it

OpenStudy (lochana):

\[\sqrt{-24} = \sqrt{-1.24} = \sqrt{i^224} = \sqrt{i^2}\sqrt{24}\]

OpenStudy (lochana):

\[\sqrt{i^2}\sqrt{24} = i\sqrt{24}\]

OpenStudy (lochana):

okay?

OpenStudy (pokiedokie):

okay i think i understand it a little better

OpenStudy (lochana):

yes. now you can do further simplifying \[i\sqrt{24} = i\sqrt{2^2.6}\]

OpenStudy (lochana):

that means \[2i\sqrt{6}\]

OpenStudy (lochana):

so \[2\sqrt{-24} + 6 = 2i\sqrt{6} + 6\]

OpenStudy (lochana):

and \[2\sqrt{-24} + 6 = 2i\sqrt{6} + \sqrt{6}\sqrt{6}\]

OpenStudy (pokiedokie):

oh wow thank you so much, i was stuck on that question for an hour...

OpenStudy (lochana):

and I did a mistake:) \[2\sqrt{-24} + 6 = \sqrt{6}(4i + \sqrt{6})\]

OpenStudy (lochana):

do it it again. you should get my last answer.

OpenStudy (lochana):

what about 2nd one?

OpenStudy (pokiedokie):

i'm not exactly sure how to do it it's confusing me a lot

OpenStudy (lochana):

you want to get removed 'i' part from denominator.

OpenStudy (lochana):

denominator is (1 + 5i)

OpenStudy (pokiedokie):

okay, and i do that by adding to both sides?? or am i incorrect

OpenStudy (lochana):

yes. you don't add things in fractions. you can't do that.

OpenStudy (lochana):

you need to multiply both sides by (1-5i)

OpenStudy (lochana):

\[\frac{(2-3i)(1-5i)}{(1+5i)(1-5i)}\]

OpenStudy (pokiedokie):

okay, and sorry that's what i meant i'm just a little stressed out so its: (2-3i)*(1+5i) _____________ (1 + 5i) (1 + 5i)

OpenStudy (pokiedokie):

whoops

OpenStudy (lochana):

it is not (1+5i). it should be (1-5i)

OpenStudy (pokiedokie):

oh okay! that makes a lot of sense

OpenStudy (lochana):

yeah. conjugate. you want to multiply by conjugate of (1+5i)

OpenStudy (lochana):

can you do it now?

OpenStudy (pokiedokie):

yes, thank you so much

OpenStudy (lochana):

you are welcome:)